OpenStudy (hba):

A possible interval of values of k for which the equation (x+k)x=-4,have two real solutions is :

OpenStudy (hba):

k>0,k>2,k>4,k<4

OpenStudy (anonymous):

(x+k)x=-4 x^2 + kx + 4 = 0

OpenStudy (anonymous):

First, rearrange the equation into the form ax^2 + bx + c = 0 after expanding the original equation.

OpenStudy (hba):

I got that already @Yahoo!

OpenStudy (hba):

what next ?

OpenStudy (anonymous):

For two real solutions b^2 - 4ac > 0

OpenStudy (anonymous):

x^2 + kx + 4 = 0 k^2 - 16 > 0

OpenStudy (hba):

k >4 < Answer

OpenStudy (anonymous):

yup..)